# biblio.bib

### v1

@@ Appendix B Circuit encoding and embedding
Starting from a quantum circuit $\mathbf{x}$, we first tokenize the gates into an integer-matrix representation (see Fig. 6), following the method of Ref. [furrutter2024quantum]. For this, each gate is assigned to a distinct token. Moreover, a sign is added to specify the connection type, e.g. control connections are implemented as negative tokens and target connections as positive ones. The resulting matrix has then dimension equal to the number of qubits $n$ times the number of gates $t$, i.e. $\mathrm{tokenize}(\mathbf{x})\in\mathbb{Z}^{n\crossproduct t}$, which is the format we use to store the circuits of the dataset (see App. C).
In this work, extending from the previous representation, we consider additional continuous parameters for the parameterized gates. We store them alongside the token matrix as a one dimensional array matching the sequence length $t$ (see Fig. 6), i.e. $\mathrm{parameters}(\mathbf{x})\in\mathbb{R}^{t}$. Importantly, the values of this array for non-parametrized gates are set to zero.
After tokenization, in order to diffuse a circuit $\mathbf{x}$ via the forward process defined in Eq. 3, we embed it into a continuous representation $[\mathbf{h}_{0},\mathbf{w}_{0}]$, which we discuss in the following Section B.1 and Section B.2.
@@ B.1 Discrete token embedding
As described in Section 3.2, we implement the token embeddings as $d_{h}$-dimensional entries $\mathbf{h}_{0}^{(i)}\in\mathbb{R}^{d_{h}}$ of a look-up table, where $i={0,\dots,N-1}$ indexes the $N$ different classes. To ensure that all embeddings are equidistant and undergo uniform mixing throughout the diffusion process, we construct them as an orthogonal basis of $\mathbb{R}^{d_{h}}$.
@@
The first point Eq. 62 addresses the known average brightness issue of DMs [zhang2024tackling, lin2024common], where they conserve the same mean as that of the starting noise latent (at $t=1$) during the whole ancestral sampling process. This property is attributed to the fact that the low frequency information of the original data $\mathbf{x}$ survives the forward diffusion process longer than high frequency details, enabling the DM to use the low frequencies as a help to lower the loss. For instance, as images have rarely zero mean, the latent initialization of ${\bf\it\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ at $t=1$ biases the DM towards very specific images, i.e. those with zero average brightness. For circuit generation, we found the low frequency information results in the circuit length to be easily predicable during training, stemming from the fact that the padding token count offsets the mean of a whole circuit embedding linearly, when not guaranteeing property Eq. 62. The second point Eq. 63 accounts for a variance preserving diffusion process [karras2022elucidating], where we require normalized data with unit variance.
Remarkably, by considering the constraints Eq. 62 and Eq. 63 for single token embeddings, all possible circuits are always embedded with zero-mean and unit-variance by construction. Note, the orthogonality together with the constraints (Eq. 62 and Eq. 63) allow a maximum of $N=d_{h}-1$ classes, as the zero-sum reduces one dimension. For the gate set of Section 4 (i.e. tokens $\left\{\mathrm{empty},\mathrm{h},\pm\mathrm{cx},\pm\mathrm{ccx},\mathrm{swap},\mathrm{rx},\mathrm{ry},\mathrm{rz},\mathrm{cp},\mathrm{padding}\right\}$) we have $N=12$ connection types in total. Therefore, we use the embedding dimension of $d_{h}=13$.
@@ B.2 Continuous parameter embedding
where both $\mathbf{v}_{1},\mathbf{v}_{2}\in\mathbb{R}^{d_{w}}$ fulfill the conditions Eq. 62 and Eq. 63. Further, we set $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ to be orthogonal, fixing $d_{w}=3$. Given a noisy embedding $\mathbf{w}_{t}$, we can decode the parameter using the estimator
$\hat{\lambda}_{t}=\frac{1}{\pi}\mathrm{arctanh2}\left(\frac{\innerproduct{\mathbf{v}_{2}}{\mathbf{w}_{t}}}{\innerproduct{\mathbf{v}_{1}}{\mathbf{w}_{t}}}\right).$ (65)
@@
In Fig. 2b we show the $\mathrm{CircularLoss}(t)$ of different noise schedules, which are defined as follows:
@@
    Linear: The linear beta schedule of DDPM [ho2020denoising]: $\beta_{t}^{w}=(1-t)\beta_{0}+t\beta_{1}$, with $\beta_{0}^{w}=10^{-4}$ and $\beta_{1}^{w}=0.02$.
@@
    CosineAlpha: The cosine schedule [nichol2021improved]: $\bar{\alpha}_{t}^{w}=\cos(t\pi/2)$.
@@
    CosineAlpha2: The squared version of the cosine schedule: $\bar{\alpha}_{t}^{w}=\cos(t\pi/2)^{2}$.
@@
In Section 4, we use for all experiments the CosineAlpha2 noise schedule for the continuous mode (see Fig. 2b and c).
Appendix C Dataset details

### v2

@@ Appendix B Circuit encoding and embedding
Starting from a quantum circuit $\mathbf{x}$, we first tokenize the gates into an integer-matrix representation (see Fig.˜9), following the method of Ref. [30]. For this, each gate is assigned to a distinct token. Moreover, a sign is added to specify the connection type, e.g. control connections are implemented as negative tokens and target connections as positive ones. The resulting matrix has then dimension equal to the number of qubits $n$ times the number of gates $t$, i.e. $\mathrm{tokenize}(\mathbf{x})\in\mathbb{Z}^{n\crossproduct t}$, which is the format we use to store the circuits of the dataset (see App.˜D).
In this work, extending from the previous representation, we consider additional continuous parameters for the parameterized gates. We store them alongside the token matrix as a one dimensional array matching the sequence length $t$ (see Fig.˜9), i.e. $\mathrm{parameters}(\mathbf{x})\in\mathbb{R}^{t}$. Importantly, the values of this array for non-parametrized gates are set to zero.
After tokenization, in order to diffuse a circuit $\mathbf{x}$ via the forward process defined in Eq.˜3, we embed it into a continuous representation $\mathbf{h}_{0},\mathbf{w}_{0}]$, which we discuss in the following Section˜B.1 and Section˜B.2.
@@ B.1 Discrete token embedding
As described in Section˜III.2, we implement the token embeddings as $d_{h}$-dimensional entries $\mathbf{h}_{0}^{(i)}\in\mathbb{R}^{d_{h}}$ of a look-up table, where $i={0,\dots,N-1}$ indexes the $N$ different classes. To ensure that all embeddings are equidistant and undergo uniform mixing throughout the diffusion process, we construct them as an orthogonal basis of $\mathbb{R}^{d_{h}}$.
@@
The first point Eq.˜62 addresses the known average brightness issue of DMs [72, 73], where they conserve the same mean as that of the starting noise latent (at $t=1$) during the whole ancestral sampling process. This property is attributed to the fact that the low frequency information of the original data $\mathbf{x}$ survives the forward diffusion process longer than high frequency details, enabling the DM to use the low frequencies as a help to lower the loss. For instance, as images have rarely zero mean, the latent initialization of ${\bf\it\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ at $t=1$ biases the DM towards very specific images, i.e. those with zero average brightness. For circuit generation, we found the low frequency information results in the circuit length to be easily predicable during training, stemming from the fact that the padding token count offsets the mean of a whole circuit embedding linearly, when not guaranteeing property Eq.˜62. The second point Eq.˜63 accounts for a variance preserving diffusion process [31], where we require normalized data with unit variance.
Remarkably, by considering the constraints Eq.˜62 and Eq.˜63 for single token embeddings, all possible circuits are always embedded with zero-mean and unit-variance by construction. Note, the orthogonality together with the constraints (Eq.˜62 and Eq.˜63) allow a maximum of $N=d_{h}-1$ classes, as the zero-sum reduces one dimension. For the gate set of Section˜IV (i.e. tokens $\left\{\mathrm{empty},\mathrm{h},\pm\mathrm{cx},\pm\mathrm{ccx},\mathrm{swap},\mathrm{rx},\mathrm{ry},\mathrm{rz},\mathrm{cp},\mathrm{padding}\right\}$) we have $N=12$ connection types in total. Therefore, we use the embedding dimension of $d_{h}=13$.
@@ B.2 Continuous parameter embedding
where both $\mathbf{v}_{1},\mathbf{v}_{2}\in\mathbb{R}^{d_{w}}$ fulfill the conditions Eq.˜62 and Eq.˜63. Further, we set $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ to be orthogonal, fixing $d_{w}=3$. Given a noisy embedding $\mathbf{w}_{t}$, we can decode the parameter using the estimator
[…]
@@
In Fig.˜2b we show the $\mathrm{CircularLoss}(t)$ of different noise schedules, which are defined as follows:
@@
    Linear: The linear beta schedule of DDPM [32]: $\beta_{t}^{w}=(1-t)\beta_{0}^{w}+t\beta_{1}^{w}$, with $\beta_{0}^{w}=10^{-4}$ and $\beta_{1}^{w}=0.02$.
@@
    Cosine: The cosine schedule [39]: $\bar{\alpha}_{t}^{w}=\cos(t\pi/2)$.
@@
    Cosine squared: The squared version of the cosine schedule: $\bar{\alpha}_{t}^{w}=\cos(t\pi/2)^{2}$.
@@
In Section˜IV, we use for all experiments the cosine squared noise schedule for the continuous mode (see Fig.˜2b and c).
Appendix C Additional experiments
C.1 Comparison to existing methods: QFT
Here we compare the compilers of Section˜IV.2.1 on a single unitary – the Quantum Fourier transform (QFT) unitary. We report in Table˜2 the synthesis results and the known textbook solution [60]. Here, we use default compiler settings, without restricting their runtimes. For genQC2, we generate 2048 circuits and select the circuit with minimum infidelity, as mentioned Fig.˜8. For the ansatz improvements, we use the same settings as in Section˜C.5. We find that our model and the ansatz fixes are capable of finding the exact textbook circuit, only missing some numeric precision in the optimized angles.
Table 2: Synthesis methods comparison: QFT. Infidelity and gate count of found circuits for the QFT unitary.
Method # qubits Infidelity Gate count
Textbook circuit 3 7
(Ref. [60]) 4 12
5 17
genQC2 (ours) 3 $1.4\cdot 10^{-4}$ 7
4 $3.8\cdot 10^{-3}$ 12
5 $7.8\cdot 10^{-2}$ 15
+ ansatz fixes 3 $1.3\cdot 10^{-7}$ 7
(Section˜IV.2.2) 4 $9.2\cdot 10^{-7}$ 12
5 $3.4\cdot 10^{-9}$ 17
QSD 3 $8.9\cdot 10^{-16}$ 61
(Refs. [54, 55]) 4 $3.6\cdot 10^{-15}$ 281
5 $3.0\cdot 10^{-12}$ 1540
AQC 3 $2.0\cdot 10^{-9}$ 79
(Ref. [13]) 4 $5.9\cdot 10^{-5}$ 317
5 $1.0\cdot 10^{-4}$ 1275
LEAP 3 $8.9\cdot 10^{-16}$ 70
(Ref. [11]) 4 $6.7\cdot 10^{-16}$ 130
5 $1.2\cdot 10^{-14}$ 212
C.2 Comparison to existing methods: restricted gate count
In this section, we want to compare our method to the approximate compilers of Table˜1, i.e. AQC and LEAP, with the new addition of QFAST [12]. Importantly, here we limit the amount of multi-qubit gates the comilers are allowed to use and report the results in Table˜3. As in Section˜IV.2.1, to speed up the simulations, we limit the number of circuit samples per unitary and timeout compilations which take more than one hour for one unitary. For LEAP and QFAST, we sample for 3 qubits 4 circuits per unitary, for 4 qubit 2 circuits per unitary, and for 5 qubits only one. Further, for AQC we sample for 3, 4 and 5 qubits each 128 circuits per unitary.
Table 3: Synthesis methods comparison with restricted gate counts. Reported values are averages over random 480 unitaries (from 2 to 16 gates) and 128 circuits sampled for each unitary. Listed errors are the corresponding standard deviation. We limit the number of CNOT layers for AQC, for LEAP the number of multi-qubit gates, and QFAST the number of 2-qubit gates. The lower bound of AQC refers to the theoretical lower bound of required CNOT gates [13]. (*) Some sample restrictions apply, as detailed in Section˜IV.2.1 and Section˜C.2. (**) Not reported due to lower sample count.
Method # qubits Limit # of multi-qubit gates Avg. minimum infidelity Avg. gate count Avg. runtime per sample (seconds) Avg. distinct circuits per unitary
genQC2 (ours) 3 NA $0.09\pm 0.20$ $8\pm 4$ $0.09\pm 0.00$ $78\pm 45$
4 NA $0.10\pm 0.20$ $8\pm 4$ $0.09\pm 0.00$ $74\pm 49$
5 NA $0.09\pm 0.17$ $8\pm 4$ $0.09\pm 0.00$ $71\pm 50$
+ ansatz fixes 3 NA $0.008\pm 0.037$ $9\pm 5$ $190\pm 390$ NA
(Section˜IV.2.2) 4 NA $0.015\pm 0.062$ $9\pm 5$ $330\pm 570$ NA
5 NA $0.015\pm 0.071$ $9\pm 5$ $480\pm 880$ NA
AQC* 3 1 $0.67\pm 0.20$ $13\pm 1$ $0.01\pm 0.00$ $7\pm 8$
(Ref. [13]) 2 $0.63\pm 0.17$ $18\pm 1$ $0.02\pm 0.01$ $12\pm 13$
4 $0.31\pm 0.23$ $28\pm 1$ $0.03\pm 0.01$ $32\pm 26$
(low. bound)  14 $(1.0\pm 3.8)\cdot 10^{-9}$ $79\pm 0$ $0.3\pm 0.2$ $21\pm 12$
16 $(6.1\pm 7.5)\cdot 10^{-10}$ $89\pm 0$ $0.3\pm 0.1$ $10\pm 7$
4 1 $0.80\pm 0.15$ $16\pm 1$ $0.02\pm 0.01$ $10\pm 10$
2 $0.81\pm 0.13$ $21\pm 1$ $0.03\pm 0.01$ $13\pm 15$
4 $0.73\pm 0.15$ $31\pm 1$ $0.04\pm 0.02$ $25\pm 24$
16 $0.037\pm 0.081$ $88\pm 2$ $0.3\pm 0.2$ $97\pm 33$
(low. bound)  61 $(4.3\pm 5.4)\cdot 10^{-5}$ $317\pm 0$ $5.2\pm 0.1$  ∗∗
5 1 $0.83\pm 0.18$ $19\pm 2$ $0.04\pm 0.01$ $12\pm 12$
2 $0.88\pm 0.10$ $24\pm 1$ $0.05\pm 0.02$ $13\pm 14$
4 $0.90\pm 0.06$ $34\pm 1$ $0.07\pm 0.03$ $18\pm 17$
16 $0.28\pm 0.26$ $91\pm 2$ $0.3\pm 0.2$ $104\pm 27$
(low. bound)  252 $(1.0\pm 0.3)\cdot 10^{-4}$ $1275\pm 0$ $27.7\pm 0.1$  ∗∗
LEAP* 3 1 $0.44\pm 0.27$ $27\pm 3$ $1.2\pm 0.1$  ∗∗
(Ref. [11]) 2 $0.22\pm 0.23$ $38\pm 7$ $1.9\pm 0.5$  ∗∗
4 $0.028\pm 0.065$ $53\pm 17$ $3.0\pm 9.5$  ∗∗
16 $(1.5\pm 3.8)\cdot 10^{-3}$ $60\pm 26$ $1.6\pm 0.6$  ∗∗
no limit $(0.4\pm 1.9)\cdot 10^{-3}$ $61\pm 26$ $0.8\pm 0.6$ $41\pm 39$
4 1 $0.56\pm 0.30$ $32\pm 3$ $1.6\pm 0.2$  ∗∗
2 $0.36\pm 0.30$ $43\pm 6$ $5.3\pm 3.0$  ∗∗
4 $0.14\pm 0.21$ $59\pm 16$ $6\pm 21$  ∗∗
16 $(2.5\pm 6.7)\cdot 10^{-3}$ $86\pm 50$ $25\pm 85$  ∗∗
no limit $(1.2\pm 3.2)\cdot 10^{-3}$ $85\pm 50$ $40\pm 250$  ∗∗
5 1 $0.59\pm 0.34$ $6\pm 3 3.3$$3\pm 1.1$  ∗∗
2 $0.42\pm 0.35$ $47\pm 7$ $50\pm 40$  ∗∗
4 $0.21\pm 0.29$ $63\pm 17$ $45\pm 77$  ∗∗
16 $0.012\pm 0.041$ $95\pm 61$ $240\pm 600$  ∗∗
no limit $(2.7\pm 5.3)\cdot 10^{-3}$ $92\pm 57$ $170\pm 500$  ∗∗
QFAST* 3 1 $0.65\pm 0.24$ $20\pm 2$ $1.6\pm 0.1$  ∗∗
(Ref. [12]) 4 $0.25\pm 0.21$ $47\pm 10$ $1.7\pm 0.2$  ∗∗
16 $0.046\pm 0.070$ $80\pm 30$ $2.0\pm 0.6$  ∗∗
36 $(2.7\pm 5.2)\cdot 10^{-3}$ $99\pm 45$ $2.4\pm 1.2$  ∗∗
4 1 $0.84\pm 0.20$ $20\pm 4$ $1.9\pm 0.2$  ∗∗
4 $0.65\pm 0.27$ $45\pm 10$ $2.3\pm 0.3$  ∗∗
16 $0.38\pm 0.27$ $80\pm 23$ $2.9\pm 0.8$  ∗∗
36 $0.23\pm 0.22$ $115\pm 37$ $4.0\pm 1.7$  ∗∗
5 1 $0.92\pm 0.15$ $19\pm 4$ $2.7\pm 0.5$  ∗∗
4 $0.83\pm 0.22$ $44\pm 11$ $3.5\pm 0.8$  ∗∗
16 $0.65\pm 0.30$ $80\pm 20$ $6.5\pm 2.7$  ∗∗
36 $0.54\pm 0.32$ $115\pm 33$ $13.0\pm 8.1$  ∗∗
C.3 Comparison to genQC1
We compare the method presented in this paper (genQC2) to its predecesor genQC1 [30]. To summarize, the major upgrades from genQC1 are:
  1. 1.
    New multi-modal diffusion with separated discrete and continuous modes, allowing native generation of parameterized gates jointly with their continuous control parameters, without post-gradient-based parameter optimization. (App.˜A)
  2. 2.
    Completely new transformed-based architecture (CirDiT; App.˜G).
  3. 3.
    Pre-training of a custom Unitary-CLIP (App.˜H).
  4. 4.
    Scaling to 5 qubits with up to 32 gates (App.˜D).
  5. 5.
    General improvements in learning: learned noise schedulers to better account for the variance of both modes, adjusted loss weighting with mode specific signal-to-noise ratios, improved discrete embedding choice (zero mean, unit variance) (App.˜A and App.˜B).
To directly compare the compilation accuracy of genQC1 and genQC2, we have to restrict the unitaries to the circuit limitations of genQC1, which are only 3 qubits and a maximum number of 12 discrete gates. In Fig.˜10, we show a comparison of the infidelities on the gate sets each of the two models were trained on, using, respectively, 725 and 650 unitaries. We sample 128 circuits per unitary and record the minimum infidelity. We note that the two models compile into their respective gate sets, while the target unitary gate set varies. In particular, genQC1 was trained on discrete circuits while genQC2 also consider continuous gates. Importantly, we observe that genQC2 is reaching a similar performance as genQC1 on the discrete gate set, without being trained explicitly for it. Contrary, genQC1 is not able to compile unitaries consisting of parameterized gates with low infidelity.
Figure 10: Comparison genQC 1 vs 2. Average minimum infidelities vs number of gates of the target unitaries. Error bars are the error of the means, and error of the average infidelity (value of the legend) is the standard deviation of the plotted points. The target unitaries are over 3 qubits, up to a maximum number of 12 gates, and are sampled with: a) the discrete gate set of genQC1; b) the discrete-continuous gate set of genQC2.
C.4 Post-optimization of continuous gates
We analyze now the accuracy of the continuous parameters of the generated circuits, i.e. the performance of the continuous mode. For this, we explore how much a given circuit can be improved using a standard gradient-based optimization method. We then compare these optimized results with those obtained from optimizing the same circuits whose continuous parameters were uniformly sampled.
For the results shown in Fig.˜11, we run the optimizer on generated circuits for 3 to 5 qubits, each 256 unitaries, uniformly consisting of 2 to 32 gates. We consider 128 generated circuits per unitary. For the gradient-based optimization of the continuous parameters, we use the backpropagation method, a learning rate (lr) of $0.1$ with a cosine annealing lr-schedule and a maximum of 100 update steps. To speed up the optimization, we stop early if the infidelity does not change anymore than $10^{-12}$.
We see in Fig.˜11 that the predicted initialization of the generated circuits has a notable lower infidelity than random starting points. We also observe a faster convergence of the generated parameters, accelerating the optimization and reducing the number of update steps required. The fact that the optimizer does not reach zero is attributed to the erroneous placement of gates in the generated ansätze. We discuss possible fixes for this in Section˜IV.2.2.
Figure 11: Post-optimization. Gradient-based optimization of the continuous parameters of the generated circuits, starting with the parameters predicted by the model or randomly initialized parameters. Shown are in a) the average mean infidelities, and in b) the average minimum infidelities. Shaded areas are the standard deviations. Details in Section˜C.4.
C.5 Tree search details
We provide here additional details on the tree search method presented in Section˜IV.2.2. We run the search on generated circuits for 3 to 5 qubits, each 480 unitaries, uniformly consisting of 2 to 16 gates. Detailed results are presented in Table˜1 and Fig.˜5. In addition, in Fig.˜12, we show the trajectories of the circuit improvements.
We perform a top-$k$ greedy expansion policy, where we set $k=15$. For this, we only expand the $k$-lowest valued nodes, measured by the infidelities, at a given depth. The available actions are $\mathrm{delete}$, $\mathrm{insert}$ or $\mathrm{replace}$ a single gate. The gate set is the same as the model is using: $\left\{\mathrm{h},\mathrm{cx},\mathrm{ccx},\mathrm{swap},\mathrm{rx},\mathrm{ry},\mathrm{rz},\mathrm{cp}\right\}$. A node expansion then considers all the possible combinations of these action with all possible gate placement possibilities. After each action, we additional perform a gradient-based optimization of the continuous parameters, before recording the new node infidelity values. To speed up the tree-search we stop expanding nodes which have an infidelity under a given $\epsilon$-threshold. We generally use $\epsilon=10^{-6}$, reflected in Fig.˜5c and Fig.˜12.
For the gradient-based optimization of the continuous parameters, we use the backpropagation method, a learning rate (lr) of $0.1$ with a cosine annealing lr-schedule and a maximum of 1000 update steps. To speed up the optimization, we stop early if the infidelity does not change anymore than $10^{-8}$.
Figure 12: Tree search paths. Infidelity trajectories of all considered circuits after every action is applied. The dotted black line represents the expansion threshold, where each trajectory stops once it passed this threshold (see Section˜C.5). Generated circuits and their optimizations are the same as in Fig.˜5.
C.5.1 What gates and angles are added?
To understand better what kind of fixes the tree-search found (see Section˜IV.2.2), we now look at what gates are inserted and what angle distribution they have after optimization. We present in Fig.˜13ab the distribution of new gates which were added by the insert or replace action of the tree-search. We see most of the time a continuous gate is added, which we attribute to the fact that enlarging the ansatz, i.e. making it more expressive, is often the best action that a greedy policy can do.
Recall, the parametrized gates are defined as [48]:
$\mathrm{rx}(\theta)$ $=\begin{pmatrix}\cos(\theta/2)&-i\sin(\theta/2)\\ -i\sin(\theta/2)&\cos(\theta/2)\end{pmatrix}$$\quad$ $\mathrm{ry}(\theta)$ $=\begin{pmatrix}\cos(\theta/2)&-\sin(\theta/2)\\ \sin(\theta/2)&\cos(\theta/2)\end{pmatrix}$,
$\mathrm{rz}(\theta)$ $=\begin{pmatrix}\exp(-i\theta/2)&\\ &\exp(i\theta/2)\end{pmatrix}$, $\mathrm{cp}(\theta)$ $=\begin{pmatrix}1&&&\\ &1&&\\ &&1&\\ &&&\exp(i\theta)\end{pmatrix}$. (67)
Note that the $\mathrm{rx}$, $\mathrm{ry}$ and $\mathrm{rz}$ gates are $4\pi$ periodic, and the $\mathrm{cp}$-gate has a period of $2\pi$. In Fig.˜13c-f, we show the angle distributions of all the added gates after the gradient-based optimization is done. For all of the gates, we find two main peaks, one around zero and one around $2\pi$. First, we see for $\theta=0$ that $\mathrm{rx}(0)=\mathrm{ry}(0)=\mathrm{rz}(0)=\mathrm{cp}(0)=\mathbf{I}$, and secondly for $\theta=2\pi$ we have $\mathrm{rx}(2\pi)=\mathrm{ry}(2\pi)=\mathrm{rz}(2\pi)=-\mathbf{I}$. Notably, both $\pm\mathbf{I}$ have no effect on the infidelity of a circuit, as they are just global phases (see Eq.˜1), effectively turning off these added gates. The distribution around these peaks, in Fig.˜13c-f, corresponds to the actual corrections the optimizer makes to the circuit ansatz. We see they are relatively close to their peaks, making the corrections rather small perturbations around the identity.
Figure 13: Tree-search fixes. Generated circuits and their optimizations are the same as in Fig.˜5. a, b) Distribution of new gates, added by the insert and replace actions. c, d, e, f) Angle distribution of the new gates of all actions, after the corresponding optimization.
C.6 Ablation study
We list here additional details on the ablation studies of Section˜IV.2.3.
Discrete noise schedule and weighting
In Table˜4, we fix the continuous noise schedule (i.e. cosine squared) while varying only the discrete schedule and switching on our loss weighting. Results are the averages of the models presented in Fig.˜6a, which are trained for 250k update steps.
Table 4: Ablation of discrete noise schedules and loss weighting. Results are averages across 3 different model sizes and for 3 to 5 qubits, more details in Section˜C.6 and Fig.˜6a. Reported errors are the error of the means. Note that the reported numbers are averages across different model sizes, showing the average influence of our improvements and not a single model performance.
Discrete noise schedule Weighting $\omega_{h}(t),\omega_{w}(t)$ Avg. mean infid. $\downarrow$ Avg. min infid. $\downarrow$
DDPM linear (Ref. [32]) constant $0.780\pm 0.004$ $0.472\pm 0.006$
Cosine (Ref. [39]) constant $0.772\pm 0.004$ $0.463\pm 0.006$
Linear learned (Sec. III.3) constant $0.709\pm 0.005$ $0.380\pm 0.006$
Linear learned (Sec. III.3) adjusted (Sec. III.2 + III.3) $\textbf{0.672}\pm 0.005$ $\textbf{0.358}\pm 0.005$
Continuous noise schedule
In Table˜5, we fix the discrete schedule (i.e. learned linear) and loss weighting (i.e. adjusted) while varying only the continuous noise schedule. For this ablation, we train models with 32M parameters for 250k update steps. We find no significant difference and use for all of our experiments the cosine squared schedule.
Table 5: Ablation of continuous noise schedules. Results are averages for 3 to 5 qubits for a single model for each setting, more details in Section˜C.6. Reported errors are the error of the means.
Continuous noise schedule Avg. mean infid. $\downarrow$ Avg. min infid. $\downarrow$
DDPM linear (Ref. [32]) $0.735\pm 0.008$ $0.418\pm 0.010$
Cosine (Ref. [39]) $0.733\pm 0.008$ $0.410\pm 0.010$
Cosine squared (see Section˜B.2) $0.733\pm 0.008$ $0.407\pm 0.010$
C.7 Test for overfitting
To test for potential overfitting of our model, we tracked training and validation losses during training of our models. We did not find any stochastic significant difference between the two losses, indicating no sever memorization. Additionally, in Table˜6, we compare the synthesis performance on unitaries of the train and test set, confirming good generalization.
Table 6: Test set vs. train set performance. Reported values are averages over random 480 unitaries (from 2 to 16 gates), from the train or test set, and 128 circuits sampled for each unitary. Listed errors are the corresponding standard deviation.
Method # qubits Avg. minimum infidelity Avg. gate count Avg. runtime per sample (seconds) Avg. distinct circuits per unitary
genQC2 3 $0.08\pm 0.17$ $8\pm 4$ $0.09\pm 0.00$ $76\pm 46$
(train set) 4 $0.10\pm 0.21$ $8\pm 4$ $0.09\pm 0.00$ $77\pm 48$
5 $0.09\pm 0.18$ $8\pm 4$ $0.09\pm 0.00$ $75\pm 50$
genQC2 3 $0.09\pm 0.20$ $8\pm 4$ $0.09\pm 0.00$ $78\pm 45$
(test set) 4 $0.10\pm 0.20$ $8\pm 4$ $0.09\pm 0.00$ $74\pm 49$
5 $0.09\pm 0.17$ $8\pm 4$ $0.09\pm 0.00$ $71\pm 50$
Appendix D Dataset details